# Bayesian and Multi-Objective Decision Support for Real-Time Cyber-Physical Incident Mitigation

### v1

@@ 4.2 Railway CBTC System Attack
Figure 8: Bayesian Network representation of a hypothetical cyber-physical attack on a railway CBTC system. Nodes shaded in yellow denote CPS assets; blue nodes represent vulnerabilities; and purple nodes correspond to hazards. The grey-shaded node indicates attack feasibility, while the red-shaded node reflects the attacker’s objective, which is to disrupt train operation.
As the attack scenario is hypothetical, the vulnerabilities in this case study are not linked to any known CVEs. To address this limitation, CVSS 3.1 vectors were constructed by referencing analogous vulnerabilities catalogued in the NVD, which bear semantic and structural alignment with the hypothetical vulnerabilities. Initial exposure probability estimates were heuristically derived using Equation 2, and subsequently refined using the Bayesian latent variable approach outlined in Equation 7.
@@
Table 5 presents a side-by-side comparison of the conditional probability distributions for exposure and impact associated with the attacker’s goal:
H3_Disrupt_Train_Operation, derived from both CVSS-based and EPSS-based prior distributions. Notably, the EPSS-based result closely mirrors the CVSS-based result, suggesting that both scoring systems yield similar risk estimates for this attack scenario.
The CVSS-based result indicates a posterior probability of attack success of $P(E)=0.1256$ and a conditional probability of severe impact of $P(I)=1.0000$, yielding a composite risk score of $R=P(E)\times P(I)=12.56\%$. While this score is relatively modest, the analysis reveals full certainty of a critical system impact if the attack path is realised. This suggests a need for further consideration of preventive strategies tailored to this attack scenario.
@@
CVSS-based priors and EPSS-based priors, to evaluate differences in outcomes. In both cases, the Pareto fronts (Figure 9) reveal a consistent linear relationship between the probability of severe impact and CPS availability at elevated impact levels. Dense solution clusters are also observed near $\mathit{Likelihood}=1$ and $\mathit{Impact}=1$, indicating strong model confidence that attacks are likely to succeed under most mitigation configurations. This suggests that the optimisation process prioritises improvements in CPS availability, rather than prevent compromise altogether.
@@
These observations provide useful guidance for CPS practitioners. Importantly, the results suggest flexibility in using EPSS and CVSS in scenarios where none of the vulnerabilities are linked to known CVEs. Frequency analysis can further support comparative evaluation of CVSS- and EPSS-based results, enabling selection of more targeted mitigation strategies tailored to meet tight resource and real-time constraints in CPS incidents.
@@ 5 Discussion and future work
All three case studies, including the solar PV inverter running example in Section 3, demonstrate the applicability and utility of our proposed framework in supporting real-time, adaptive decision-making needs for real-world attacks. As new information becomes available, elements such as CPS assets, probabilistic estimates of attack likelihood, impact severity, system availability, and candidate countermeasure profiles, comprising mitigation probabilities, are continuously updated via a feedback loop driven by real-time monitoring and computationally efficient multi-objective optimisation. Future work could further investigate how CVSS-based priors, both calibrated and non-calibrated, and probabilistic EPSS-based priors affect multi-objective optimisation outcomes in CPS contexts.
@@
The optimisation experiments conducted in this research underscore the critical role of concurrency in enhancing computational efficiency and responsiveness for multi-objective decision-making in CPS. Using the Python concurrent.futures module, we achieved substantial reductions in cumulative optimisation time (Table 8).
@@
To further strengthen real-time decision-making capabilities in CPS, future work should explore several complementary directions that address the key challenges identified in our research.
First, our research highlighted the promise of frequency-based heuristics as a supplementary analytical layer beyond individual optimisation outcomes. However, their empirical nature prompts a critical question: To what extent can such heuristically derived recommendations be relied upon in real-time, safety-critical CPS environments? Addressing this question presents a valuable avenue for future work, particularly through comparative evaluation with empirical data.
Secondly, the stability of Pareto fronts across repeated optimisation trials remains central to ensuring decision robustness. Low trial counts were shown to produce high variance in outcomes, which may undermine the trustworthiness of mitigation recommendations. Future investigations should therefore develop convergence diagnostics to establish sufficiency thresholds for trial configurations, and to measure the robustness of optimisation solutions.
Thirdly, the incorporation of hybrid EPSS–CVSS exposure probability assessments, as proposed in our research, presents a promising direction for improving vulnerability assessment. Such hybrid models are particularly well suited to guiding mitigation strategies that must adapt in real time to both known vulnerabilities and evolving threat intelligence.
@@
Collectively, these directions aim to further improve a real-time, adaptive decision support framework for mitigating cyber incidents in CPS.
@@ 6 Related work
BNs for Decision-Making
BNs have proven effective in synthesising heterogeneous data for cybersecurity modelling, particularly where historical attack data is scarce [8, 20]. Prior studies [41, 33, 39, 23] demonstrate their value for dynamic decision-making. Bhosale et al. [5, 6] further integrate hazard, vulnerability, and asset connectivity into Bayesian Belief Networks (BBNs), supported by AutomationML for real-time data exchange.
Nonetheless, existing frameworks do not account for the unintended consequences of CPS mitigations, such as asset degradation or failure from software patching, and do not account for legacy system constraints and resource limitations. Their reliance on CVSS v3.1 also limits adaptability to evolving threats, where EPSS [21] offers complementary probabilistic insight. Crucially, these models lack feedback mechanisms to iteratively refine risk postures in response to new threats.
Our research addresses these gaps by proposing a consequence-aware, iterative framework that couples hybrid CVSS-EPSS exposure probability estimation with adaptive decision-support for CPS security.
@@
In addition to modelling attack propagation, effective CPS decision-making must account for the consequences of cyber-physical attacks. Kim et al. [24] incorporate a consequence layer within BN models to quantify impact propagation in exploited ICSs, detailing functional interdependencies and infrastructure-wide effects. ACTISM  [19] similarly adopts a consequence-driven approach for dynamic, iterative security modelling in automotive systems. In parallel, Amro et al. [3] propose a consequence-driven framework rooted in FMECA, well-aligned with CPS environments.
Although these approaches support iterative risk assessment, their transformation into adaptive decision-support tools hinges on modelling realistic, multi-stage cyber-physical attacks. To meet this challenge, our research integrates system-level and functional dependencies into BN structures informed by real-world scenarios, and explicitly models uncertainty within exposure computations. Together, these enhancements extend consequence-aware modelling toward responsive CPS decision-making under evolving threat landscapes.
DSS for CPS
Research has explored diverse methodologies for real-time CPS decision-making under uncertainty. Zaman et al. [44] propose a Markov Decision Process (MDP) framework for optimising responses to hazards such as fire outbreaks and power outages using CPS data. Javornik et al. [22] introduce a mission-centric decision support system to guide cybersecurity experts in selecting resilient IT configurations in stakeholder-driven contexts. While both frameworks offer structured resilience assessments, they do not account for CPS-specific constraints, adversary tactics, and cascading cyber-physical risks.
To address these limitations, our research employs BN structures with nodes representing CPS assets, vulnerabilities, and hazards, along with their interdependencies. This approach facilitates dynamic modelling of risk propagation and supports adaptive decision-making that reflects evolving threat scenarios and system-level complexity.
@@
Decision-making in CPS cyber incident response must balance cybersecurity, safety, and operational continuity. Li et al. [29] propose a multi-objective optimisation framework for intrusion response in ICSs, prioritising Pareto-optimal security strategies using a distance-based metric. However, it does not account for unknown vulnerabilities or emerging attack vectors, limiting responsiveness in dynamic environments. Zebrowski et al. [45] present a similar optimisation-based approach using Bayesian Networks to assess cascading cyberattack impacts and minimise catastrophic consequences. Yet, their framework assumes static cybersecurity configurations and lacks adaptability to evolving threats.
To address these limitations, our research uses a Bayesian confidence calibration technique to model uncertainty from unknown or emerging attack vectors. It also implements an iterative countermeasure portfolio refinement cycle, guided by feedback from previously evaluated mitigation strategies. These innovations support the development of adaptive, context-aware CPS security strategies tailored for real-time decision-making under uncertainty.
@@
Several studies have examined railway signalling security, with varying emphasis on CBTC systems. Xu et al. [42] present a simulation platform architecture grounded in real-world CBTC characteristics. In contrast, Schlehuber et al. [35] focus on interlocking systems without addressing CBTC explicitly, and Unger et al. [38] survey railway attack scenarios and countermeasure maturity, yet do not concentrate on CBTC systems. Conversely, Yu et al. [43] offer a comprehensive review of CBTC architecture, limitations, cybersecurity threats, and protection strategies.
Regarding systemic vulnerabilities, Farooq et al. [14] analyse over-the-air communications in CBTC, identifying interference susceptibility, particularly in Wi-Fi technologies. Oliveira et al. [32] echo this concern, using Failure Mode and Effects Analysis (FMEA) to uncover risks such as message spoofing and MitM attacks that compromise communication integrity. Similarly, Soderi et al. [36] outline wireless attack scenarios with demonstrated impacts on railway safety. These findings directly inform the attack modelling applied in the CBTC case study presented in this paper.
@@ 7 Conclusion
We presented a real-time, adaptive decision support framework for mitigating cyber incidents in CPS environments. Its development is timely, given the growing reliance on CPS within critical infrastructure, and the rapidly evolving cyber threat landscape.
A key contribution of this work is the use of BN models to represent attack scenarios, enriched with CPS system and vulnerability metadata to facilitate probabilistic reasoning and risk-informed recommendations for real-time incident response. The models are encoded in a DSL format to enhance computational efficiency under operational constraints, and to support dynamic updates for adaptive decision support.
We also introduced a novel decision-support methodology centred on countermeasure portfolios derived through multi-objective optimisation and frequency-based heuristic analysis. Mitigation probabilities are treated as parameters within a structured optimisation search space, allowing the generation of Pareto-optimal portfolios that balance system availability, attack success likelihood, and impact severity. frequency-based heuristic analysis across multiple optimisation runs is further employed to identify consistently effective mitigation nodes, thereby guiding prioritisation decisions under time and resource constraints.
Additionally, we used a hybrid exposure probability estimation technique that integrates EPSS and CVSS scores via Bayesian confidence calibration. This approach yields robust and realistic exposure recommendations under epistemic uncertainty arising from incomplete or heterogeneous vulnerability data.
The framework was evaluated across three representative CPS attack scenarios, encompassing both multi-path and multi-agent threats. Collectively, these evaluations demonstrate the framework’s utility in supporting real-time, adaptive decision-making for cyber incident mitigation in CPS environments.
@@
The authors declare that there are no conflicts of interest that could have influenced the conduct or outcomes of this research. Furthermore, this research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
@@
The models and source code developed as part of this research are publicly available via the project’s GitHub repository ([18]).

### v2

@@ 4.2 Railway CBTC System Attack
Figure 8: Bayesian Network of a hypothetical cyber-physical attack on a railway CBTC system. Node colours denote assets (orange), vulnerabilities (green), hazards (yellow), attack feasibility (cyan), and attacker goal (red). The graph shows how vulnerability and attack techniques cascade to CBTC subsystems, supporting probabilistic risk assessment. A high-resolution version is available in the project’s GitHub repository ([20]).
As the attack scenario is hypothetical, the vulnerabilities in this case study are not linked to any known CVE entries. To address this limitation, CVSS 3.1 vectors were constructed by referencing analogous vulnerabilities catalogued in the NVD, which bear semantic and structural alignment with the hypothetical vulnerabilities. Initial exposure probability estimates were heuristically derived using Equation 3, and subsequently refined using the Bayesian latent variable approach outlined in Equation 5.
To validate the robustness of these estimates, we perform a comparative analysis against an EPSS-based prior. Although EPSS scores are natively unavailable for non-CVE vulnerabilities, this comparison serves as a sensitivity check to determine if the exposure probabilities remain consistent when calibrated against empirical exploitation trends. By contrasting the CVSS-based estimates with a synthetic EPSS-based prior, we can assess whether the risk assessment remains stable across different probabilistic assumptions regarding the threat landscape.
@@
Table 7 presents a side-by-side comparison of the conditional probability distributions for exposure and impact associated with the primary attacker goal: H3_Disrupt_Train_Operation. The CVSS-based results indicate a posterior probability of attack success of $P(E)=0.1256$ (using an AF modifier of $\phi=1.0$) and a conditional probability of severe impact of $P(I)=1.0000$, yielding a risk score of $R=P(E)\times P(I)=12.56\%$. Although this score is relatively low, the analysis reveals absolute certainty of a critical system impact if the attack path is realised. This findings highlights a significant vulnerability within the system architecture. While the likelihood of exploitation is limited, the lack of resilience against a successful breach ensures a catastrophic outcome. Consequently, the results suggest that preventive strategies must be prioritised to reduce the probability of entry, as the current system configuration offers no mitigation for the impact itself once the goal is achieved.
@@
CVSS-based priors and EPSS-based priors to evaluate differences in outcomes. In both cases, the derivation of the Pareto fronts (Figure 9) illustrates a trade-off between exposure, impact, and availability. The results reveal a consistent relationship where the likelihood of severe impact remains elevated across all levels of attack success. Dense solution clusters are observed near $\textit{Likelihood}\approx 1$, indicating strong model confidence that attacks are likely to succeed under most mitigation configurations. Furthermore, the analysis demonstrates that the probability of severe impact remains significant, typically ranging between 0.55 and 0.75. This suggests that the optimisation process prioritises incremental improvements to maximise system availability, which remains constrained within the 0 to 0.4 range, rather than seeking to prevent compromise or fully suppress the resulting impact altogether.
@@
4.3 Trade-offs and Practitioner Guidance
In interpreting the results of our three case studies, it is important to distinguish scenario-specific behaviours from properties of the framework itself. Patterns such as the shapes of Pareto fronts, the prominence of specific vulnerabilities, or clustering effects in the BlackEnergy and CBTC experiments stem from the particular architectures, threat models, and parameter settings of those scenarios. In contrast, the ability to generate Pareto-optimal countermeasure portfolios, quantify trade-offs between exposure, impact, and availability, and rank mitigation options via frequency-based heuristics are general characteristics of our proposed method that should extend to a broad range of CI deployments.
These findings provide practical guidance for CI practitioners regarding the navigation of security trade-offs. The results indicate that complementary vulnerability scoring metrics such as CVSS, EPSS, LEV, and KEV can be used flexibly even when vulnerabilities lack CVE identifiers. Specifically, the integration of LEV and KEV metrics ensures that confirmed exploitation evidence and local environmental factors take precedence in the prioritisation process. This multidimensional approach, supported by frequency analysis for the comparative evaluation of results, enables more targeted mitigation strategies under tight resource and real-time constraints.
A central insight from our evaluation is that incident mitigation in CI inherently involves a three-way trade-off between exposure, impact, and system availability. Across all case studies, the optimisation results reveal regions where reducing the likelihood of successful attack or severe impact begins to incur disproportionately high availability costs, as well as regions where additional mitigation yields diminishing returns in risk reduction. In practice, operators must therefore balance tolerable residual exposure and impact against the need to keep CI services online, rather than pursuing maximal mitigation in isolation.
@@ 5 Discussion and Future Work
All three case studies, including the solar PV inverter running example in Section 3, demonstrate the applicability and utility of our proposed framework in supporting real-time, adaptive decision-making needs for real-world attacks. As new information becomes available, elements such as CPS assets, probabilistic estimates of attack likelihood, impact severity, system availability, and candidate countermeasure profiles, comprising mitigation probabilities, are continuously updated via a feedback loop driven by real-time monitoring and computationally efficient multi-objective optimisation.
From a deployment perspective, the framework presupposes access to architectural knowledge, vulnerability metadata such as CVEs, and attack telemetry to update evidence in the BN. It is designed to integrate as an analytical back-end that can be invoked from existing Security Operations Centre (SOC) workflows. In particular, the framework can serve as a complementary decision-support service that consumes alerts and configuration data, rather than replacing existing SOC tooling. Our experiments in this work demonstrate that multi-objective optimisation can be parallelised to ensure runtimes compatible with operational time-frames. However, practical deployments will require engineering trade-offs between model granularity, update frequency, and available compute resources to ensure acceptable overhead in real-time or near–real-time settings.
@@
The optimisation experiments conducted in this research underscore the critical role of concurrency in enhancing computational efficiency and responsiveness for multi-objective decision-making in CI. Using the Python concurrent.futures module, we achieved substantial reductions in cumulative optimisation time (Table 10).
@@
To further strengthen real-time decision-making capabilities in CI environments, future work should explore several complementary directions that address the key challenges identified in our research.
First, our research highlighted the promise of frequency-based heuristics as a supplementary analytical layer beyond individual optimisation outcomes. However, their empirical nature prompts a critical question: To what extent can such heuristically derived recommendations be relied upon in real-time, safety-critical CI environments? Addressing this question presents a valuable avenue for future work, particularly through comparative evaluation with empirical data. Future work could also extend optimisation to explicitly account for resilience alongside system availability.
Second, the stability of Pareto fronts across repeated optimisation trials remains central to ensuring decision robustness. Low trial counts were shown to produce high variance in outcomes, which may undermine the trustworthiness of mitigation recommendations. Future investigations should therefore develop convergence diagnostics to establish sufficiency thresholds for trial configurations, and to measure the robustness of optimisation solutions.
Third, the incorporation of hybrid confidence-calibrated exposure probability assessments presents a promising direction for performing vulnerability assessment in practical deployments. Such hybrid approaches are particularly relevant in guiding real-world incident mitigation strategies that must adapt in real time to evolving attacks and dynamic threat intelligence. Furthermore, the approach is inherently extensible, facilitating the integration of diverse vulnerability scoring systems such as CVSS, EPSS, LEV, and KEV to refine the quantification of exposure estimates. Future work could further leverage this multidimensional approach to refine decision support, enabling a more comprehensive and data-informed basis for prioritising countermeasures as new threat intelligence becomes available.
@@
Collectively, these directions aim to further improve a real-time, adaptive decision support framework for mitigating cyber incidents in CI.
@@ 6 Related Work
Probabilistic Modelling in Cybersecurity
BNs are highly effective in synthesising heterogeneous data for cybersecurity modelling, particularly when historical attack data is sparse [8, 19]. Prior studies [44, 34, 42, 23] underscore their utility in dynamic decision-making. Bhosale et al. [5, 6] further integrated hazard, vulnerability, and asset connectivity into Bayesian Belief Networks (BBNs) using AutomationML for real-time data exchange. However, existing frameworks often overlook the operational trade-offs of security interventions, such as asset degradation or failure from software patching, and do not account for legacy system constraints and resource limitations. Furthermore, a heavy reliance on CVSS v3.1 scores limits adaptability to shifting threat landscapes where probabilistic insights from EPSS [21] are required. Our research addresses these gaps by proposing a consequence-aware, iterative framework that couples confidence-calibrated exposure probability with adaptive decision-support.
@@
Beyond modelling attack propagation, effective CI decision-making must account for the consequences of cyber-
physical incidents. Kim et al. [24] incorporate a consequence layer within BN models to quantify impact propagation in exploited ICSs, detailing functional interdependencies and infrastructure-wide effects. ACTISM  [18] similarly adopts a consequence-driven approach for dynamic, iterative security modelling in automotive systems. In parallel, Amro et al. [3] propose a consequence-driven framework rooted in FMECA that aligns well with CI environments. Although these approaches support iterative assessment, their transformation into adaptive tools hinges on modelling realistic, multi-stage attacks. Our research integrates system-level functional dependencies into BN structures informed by empirical failure modes, explicitly modelling uncertainty within exposure computations to enable responsive decision-making.
Decision Support for CI Incidents
Research has explored diverse methodologies for real-time CPS decision-making under uncertainty. Zaman et al. [47] propose a Markov Decision Process (MDP) framework for optimising responses to hazards such as fire and power outages. Javornik et al. [22] introduce a mission-centric DSS to guide cybersecurity experts in selecting resilient IT configurations. While providing structured resilience assessments, these frameworks often lack specific focus on the adversarial tactics and non-linear cascading risks inherent in CI incidents. To address these limitations, our research employs BN structures where nodes represent CPS assets, vulnerabilities, and hazards. This approach facilitates dynamic modelling of risk propagation and supports adaptive decision-making that reflects evolving threat scenarios and system complexity.
@@
Decision-making in CI incident response must balance cybersecurity, safety, and operational resilience. Rehak et al. [35] introduce a methodology for assessing CI resilience by evaluating robustness, recovery, and adaptive capacity. However, their model is essentially static and focused on pre-incident conditions rather than real-time response. Li et al. [29] propose a multi-objective optimisation framework for intrusion response in ICSs, prioritising Pareto-optimal strategies, yet it does not account for unknown vulnerabilities or emerging vectors. Similarly, Zebrowski et al. [48] use BNs to assess cascading impacts but assume static configurations. While Shimizu et al. [38] proposed a vulnerability management chaining framework integrating CVSS, EPSS, and KEV metrics, their work is tailored to enterprise IT and lacks the multi-objective weighting necessary to balance the high-availability requirements of CI environments. Our research addresses these limitations by using a Bayesian confidence calibration technique to model emerging threat uncertainty and implementing an iterative multi-objective optimisation cycle to tailor strategies for real-time CI incident mitigation.
@@
Several studies have examined the security of railway signalling, with varying emphasis on CBTC technologies. Xu et al. [45] present a simulation platform grounded in real-world CBTC characteristics, while Yu et al. [46] offer a comprehensive review of the architecture, cybersecurity threats, and protection strategies specific to these systems. In contrast, other research focuses on broader or legacy signalling components. For instance, Schlehuber et al. [37] examine interlocking systems without addressing CBTC explicitly, and Unger et al. [41] survey railway attack scenarios and countermeasure maturity across the sector more generally. Regarding systemic vulnerabilities, research has highlighted the risks inherent in the transition to wireless-based control. Farooq et al. [15] analyse over-the-air communications in CBTC, identifying interference susceptibility in Wi-Fi-based implementations. Oliveira et al. [33] expand on this by using Failure Mode and Effects Analysis (FMEA) to uncover risks such as message spoofing and Man-in-the-Middle (MitM) attacks that compromise communication integrity. Similarly, Soderi et al. [39] outline wireless attack scenarios with demonstrated impacts on railway safety. These findings directly inform the adversarial modelling applied in our CBTC case study, ensuring the evaluation reflects the unique cyber-physical risks inherent in modern automated rail operations.
@@ 7 Conclusion
This paper presented a real-time, adaptive decision support framework designed to help practitioners mitigate cyber-physical incidents within CI environments. The development of such a system is increasingly vital as CPSs become more deeply integrated into CIs while facing a rapidly diversifying threat landscape. A primary contribution of this work is the application of BN models to represent complex attack scenarios, enriched with system-specific and vulnerability metadata. By encoding these models in a DSL format, the framework achieves the computational efficiency required for operational time-frames, allowing for dynamic updates that support adaptive decision-making as an incident unfolds.
The framework further introduces a novel decision-support methodology for countermeasure selection based on multi-objective optimisation and frequency-based heuristic analysis. By treating mitigation probabilities as parameters within a structured search space, the system generates Pareto-optimal portfolios that balance the critical requirements of security, safety, and operational resilience. The use of frequency-based heuristics across multiple optimisation runs provides a practical bridge between complex model outputs and actionable guidance, enabling operators to identify the most robust and consistent mitigation strategies under severe resource constraints.
To address the inherent uncertainty of cyber-physical vulnerabilities, the framework incorporates a hybrid exposure probability estimation technique. This method integrates complementary vulnerability scoring metrics through a Bayesian confidence calibration process, yielding robust exposure recommendations even when faced with the incomplete or heterogeneous metadata typical of legacy CI assets.
The practical utility of this applied research was validated through evaluation across three representative scenarios: the 2015 Ukrainian power grid (BlackEnergy) attack, a solar PV inverter system and a railway signalling network. These case studies demonstrate the framework’s versatility in handling diverse multi-path and multi-agent threats within the CI domain. Finally, to support transparency and facilitate further development by both practitioners and researchers, the models, source code and reference data have been made publicly available [20]. This open approach ensures that the findings can be replicated and extended to enhance the protection of the essential systems that underpin modern society.
@@
The authors declare that there are no conflicts of interest that could have influenced the conduct or outcomes of this research. Furthermore, this research did not receive any specific grant from funding agencies in the public, commercial or not-for-profit sectors.
@@
The models and source code from this research, along with high-resolution versions of the paper’s figures are publicly available in the project’s GitHub repository ([20]).